# T-LoRA: Single Image Diffusion Model Customization Without Overfitting

### v1

@@ Abstract
While diffusion model fine-tuning offers a powerful approach for customizing pre-trained models to generate specific objects, it frequently suffers from overfitting when training samples are limited, compromising both generalization capability and output diversity. This paper tackles the challenging yet most impactful task of adapting a diffusion model using just a single concept image, as single-image customization holds the greatest practical potential. We introduce T-LoRA, a Timestep-Dependent Low-Rank Adaptation framework specifically designed for diffusion model personalization. In our work we show that higher diffusion timesteps are more prone to overfitting than lower ones, necessitating a timestep-sensitive fine-tuning strategy. T-LoRA incorporates two key innovations: (1) a dynamic fine-tuning strategy that adjusts rank-constrained updates based on diffusion timesteps, and (2) a weight parametrization technique that ensures independence between adapter components through orthogonal initialization. Extensive experiments show that T-LoRA and its individual components outperform standard LoRA and other diffusion model personalization techniques. They achieve a superior balance between concept fidelity and text alignment, highlighting the potential of T-LoRA in data-limited and resource-constrained scenarios. Code is available here.
Figure 1: Compared to standard LoRA single view fine-tuning, T-LoRA reduces overfitting related to position and background, enabling more versatile and enriched generation.
@@ 1 Introduction
Recent advances in personalization of pre-trained large-scale text-to-image diffusion models DB ; TI ; CD ; svdiff have transformed content creation, garnering significant attention from both academia and industry. This task enables pre-trained models with the ability to generate highly specific and customized outputs, such as particular objects, styles, or domains. The primary objectives of such customization are twofold: (1) ensuring high-quality preservation of the target concept, and (2) achieving precise alignment between the generated image and the input text.
Fine-tuning based customization approaches have proven effectiveness at producing high fidelity concept samples DB ; CD ; svdiff . However, they encounter fundamental limitations due to the restricted size of target datasets. This constraint often compromises generalization capabilities, resulting in undesirable artifacts in the generated outputs. Notably, background elements and positional biases from the training images frequently "leak" into new generations, substantially limiting the model’s flexibility and creative potential. Compared to standard full-weight fine-tuning, lightweight Low-Rank Adaptater (LoRA) lora offers significant advantages biderman2024lora ; Ryu2024githublora ; fan2023reinforcement ; smith2024continualdiffusion : it significantly reduces the number of trainable parameters, making it suitable for resource-constrained settings, and it is less prone to overfitting, thereby better preserving the original model’s generative capabilities.
Adapting a diffusion model using only a single concept image is the most extreme scenario. In this scenario the risk of overfitting is particularly pronounced and even lightweight methods ortogonal ; lora ; svdiff can exhibit susceptibility to overfitting. This is challenging yet most impactful task, on which we focus our research. A single-image customization holds the greatest practical potential as in practice, users rarely have multiple images of their concepts with varying backgrounds.
@@
Based on our analysis, we introduce T-LoRA, a Timestep-Dependent Low-Rank Adaptation framework for diffusion model customization. The core idea behind T-LoRA is to prioritize training capacity for less noisy timesteps while reducing the training signal for the noisier ones. To implement this, we propose a time-dependent masking strategy, called Vanilla T-LoRA, which restricts higher-rank LoRA components from participating in training and inference during noisier (higher) timesteps. Our further analysis reveals that standard LoRA adapters often exhibit an effective rank that is significantly smaller than the original rank hyperparameter. This property could limit the effectiveness of Vanilla T-LoRA, as masked and unmasked adapters may possess similar expressive power. To address this limitation, we introduce Ortho-LoRA, a novel LoRA initialization technique that ensures orthogonality between adapter components, thereby explicitly separating information flows across different timesteps. Extensive experiments demonstrate the effectiveness of T-LoRA and its individual components (Vanilla T-LoRA and Ortho-LoRA). Our proposed framework significantly outperforms existing lightweight fine-tuning approaches in single-image diffusion model personalization tasks in both metrics and user study. These findings establish a strong foundation for integrating temporal regularization and orthogonality into future diffusion model adaptation frameworks.
@@
  • We propose T-LoRA a novel Low-Rank Adaptation framework specifically designed to mitigate overfitting in diffusion model personalization. T-LoRA introduces an innovative rank masking strategy that balances training capacity across diffusion timesteps, applying stronger regularization to higher timesteps to mitigate overfitting.
  • We explore the concept of effective rank in LoRA matrices and propose Ortho-LoRA, a novel orthogonal weight initialization method. Ortho-LoRA improves the separation of information flows across timesteps and enhances effective rank utilization, further boosting the performance of T-LoRA.
@@ 2 Background
Diffusion models Diffusion models song2020denoising ; ho2020denoising are probabilistic generative models parametrized by neural networks and trained to approximate a data distribution by progressively denoising a Gaussian-sampled variable. Specifically, we focus on text-to-image diffusion models (ramesh2022hierarchical, ; ramesh2021zero, ; saharia2022photorealistic, ; stablediffusion, ) $\varepsilon_{\theta}$, which generate an image $x$ from a text prompt $P$. The model incorporates a text encoder $E_{T}$ to obtain a text embeddings $p=E_{T}(P)$ and a conditional diffusion model $\varepsilon_{\theta}$. Latent diffusion models Rombach_2022_CVPR further encode the image into a latent representation $z=E(x)$ via an encoder $E$ and reconstruct it with a decoder $D$, such that $x\approx D(z)$. Diffusion models $\varepsilon_{\theta}$ are trained to predict the noise contained in this noisy latent representation:
$\min_{\theta}\mathbb{E}_{p,t,z,\varepsilon}\left[\left\|\varepsilon-\varepsilon_{\theta}(t,z_{t},p)\right\|_{2}^{2}\right]$ (1)
@@
where $\varepsilon\sim N(0,I)$, $t$ is a diffusion timestep, and $z_{t}$ is a noisy latent code generated through a forward diffusion process $z_{t}=\text{ForwardDiffusion}(t,z,\varepsilon)$. During inference, a random noise $z_{T}\sim N(0,I)$ is iteratively denoised to recover $z_{0}$.
@@
LoRA Low-Rank Adaptation (LoRA) lora is a technique designed to efficiently fine-tune large pre-trained models. LoRA parametrizes original model weight matrix $W$ update as $\Delta W$ as a product of two low-rank matrices $A$ and $B$. The adapted parameters $\tilde{W}$ can be expressed as:$\tilde{W}=W+BA,A\in\mathbb{R}^{r\times m},B\in\mathbb{R}^{n\times r}$, where $r$ is the rank of the matrices $A$ and $B$ and $(n,m)$ is the dimensionality of $W$. To ensure that the model behaves identically to the pre-trained model at the beginning of training, usually $A$ is initialized normally and $B$ with zero values.
@@
Diffusion Model Customization A widely adopted approach for personalizing diffusion models is fine-tuning, which involves adjusting model weights to effectively generate specific concepts based on a limited set of concept images. This method addresses the challenge of aligning the model’s output with user-defined concepts, facilitating tailored image generation. To bind a new concept to a special text token $V^{*}$, $\varepsilon_{\theta}$ is fine-tuned on a small set of concept images =$\mathbb{C}=\{x\}_{i=1}^{N}$ with the following objective:
$\min_{\theta}\mathbb{E}_{p,t,z=\mathcal{E}(x),x\in\mathbb{C},\varepsilon}\left[\left\|\varepsilon-\varepsilon_{\theta}(t,z_{t},p)\right\|_{2}^{2}\right]$ (2)
where $p=E_{T}(P)$ is the text embeddings corresponding to the prompt $P=$a photo of a V*". Several strategies have been introduced to enhance this fine-tuning process, including pioneering work on pseudo-token optimization TI and diffusion model fine-tuning DB ; CD . Additionally, various lightweight parameterization techniques svdiff ; r1e ; lora ; ortogonal ; groupshouffle have emerged, aiming to reduce computational costs while preserving the model’s flexibility and mitigating overfitting.
Among these methodologies, Low-Rank Adaptation (LoRA) has gained particular attention. Despite its established nature, LoRA lora is recognized as a strong baseline in subject-driven generation for its lightweight design, high concept fidelity and strong alignment with input prompts. Furthermore, LoRA frequently serves as a foundational component in other personalization techniques, both finetuning-based disenbooth ; palp ; conceptguard and encoder-based e4t ; photomaker .
@@
Overfitting problem Many existing approaches struggle with overfitting to position and background, a challenge that worsens with fewer concept images. To address this, various techniques like concept masking elite ; dreamtuner ; realcustom , prompt augmentation styledrop ; tokenverse , regularization DB ; CD ; core ; attndreambooth , advanced attention mechanisms dreamtuner ; realcustom ; dreammatcher , and sampling strategies profusion ; photoswap ; beyondft have been explored. Unlike prior work, we investigate the root cause of overfitting, hypothesizing that fine-tuning during the noisiest timesteps of the diffusion process reinforces background elements from the training data, limiting generative flexibility. To counter this, we propose a novel strategy that reduces the concept signal during noisy timesteps in both training and inference. Additionally, we adapt the LoRA lora architecture to selectively control concept injection across timesteps, enhancing generalization and output diversity while mitigating overfitting.
@@ 3 Method
In this section, we introduce T-LoRA, a timestep-dependent low-rank adaptation framework, which aims to address an overfitting problem in diffusion model personalization task. The overview of method is shown in Figure 2(c). In the following contents, we first discuss the motivation of the proposed method in Section 3.1. Then, in Section 3.2, we introduce time-dependent LoRA rank controlling strategy, namely Vanilla T-LoRA (Figure 2(b)). After that, in Section 3.3 we analyze effective rank of trained LoRA adapters in diffusion models and propose Ortho-LoRA weight initialization, which ensures orthogonality of LoRA components and provide more effective rank control on the noisiest timesteps. Finally, in Section 3.4 we present full T-LoRA method that incorporates Vanilla T-LoRA rank controlling strategy and Ortho-LoRA weight initialization.
3.1 Motivation
Previous studies have shown that different timesteps in diffusion models play distinct roles throughout the generation process Choi2022CVPR ; Koyejo2022NIPS ; li2023autodiffusion . For example, the authors of Choi2022CVPR categorized the behavior of diffusion timesteps into three main stages: high timesteps ($t\in[800;1000]$) concentrate on forming coarse features, middle timesteps ($t\in[500;800]$) produce perceptually rich content, and the lower timesteps ($t\in[0;500]$) focus on removing residual noise. Additionally, works chang2023muse ; gao2023mdtv2 ; beyondft show that high timesteps contribute to image diversity, as inadequate representation of the prompt’s context during this stage makes it less likely to be restored in subsequent timesteps.
To investigate the role of different timesteps in fine-tuning, we conducted an experiment where we uniformly sampled timesteps from various fixed intervals (see Figure 3). The results show that fine-tuning at higher timesteps $t\in[800;1000]$ leads to rapid overfitting, causing memorization of poses and backgrounds, which limits image diversity and adherence to prompts. Importantly, these timesteps are essential for defining the overall shape and proportions of the concept. In the middle timesteps $t\in[500;800]$, the generated context became richer, and the model better reproduce fine concept details. However, we lose information related to the overall shape. For example, as illustrated in Figure 3, the boot retains fittings that correspond to the original concept, but its height is distorted, resulting in a shorter shoe. Finally, fine-tuning with lower timesteps $t\in[0;500]$ demonstrated the best alignment with text prompts and yielded the richest contextual generation. However, this approach struggled to accurately reproduce the intended concept, often generating only some textural elements and losing both shape and fine details of the object.
@@ 3.1 Motivation
Figure 2: Comparison of training methods: LoRA, the proposed Vanilla T-LoRA, and T-LoRA schemes. Modules marked with a snowflake icon represent components with frozen parameters during training, while modules marked with a fire icon indicate trainable components.
Figure 3: Motivational Experiment. The figure presents the results of fine-tuning the SD-XL model with LoRA across a fixed interval of timesteps. We find that focusing on the noisiest timesteps leads to rapid overfitting, particularly affecting the positioning and background of generated images. In contrast, shifting to earlier timesteps results in outputs that are better aligned with text prompts, yielding richer and more flexible generations. However, completely removing the influence of the noisiest timesteps ($t\in[800;1000]$) is not feasible, as they are crucial for preserving concept fidelity.
@@ 3.2 Vanilla T-LoRA
To tackle the aforementioned challenge, we propose a timestep-dependent fine-tuning strategy for diffusion models. Our approach involves parameterization that allocates fewer parameters for training at higher timesteps and more parameters at lower timesteps. Specifically, we introduce a masking mechanism to dynamically adjust the ranks of the LoRA adapters based on the timestep within the diffusion process, referred to as Vanilla T-LoRA (see Figure 2(b)):
$\centering\begin{split}&\tilde{W}_{t}=W+B_{t}A_{t}=W+BM_{t}A,\quad A\in\mathbb{R}^{r\times m},B\in\mathbb{R}^{n\times r}\\ &M_{t}=M_{r(t)}=\mathrm{diag}(\underbrace{1,1,\dots,1}_{r(t)},\underbrace{0,0,\dots,0}_{r-r(t)})\in\mathbb{R}^{r\times r}\end{split}\@add@centering$ (3)
We propose to define $r(t)$ as a linear function inversely proportional to the timesteps: $r(t)=\lfloor(r-r_{min})\cdot(T-t)/T\rfloor+r_{min}$. At the maximum timestep $T$, $r(t)$ will equal $r_{\text{min}}$, a hyperparameter of our method. Then, $r(t)$ will linearly increase, reaching a maximum value of $r$ at 00 timestep.
This rank-masking strategy enables dynamic control over the amount of information injected at different timesteps during both adapter training and inference. Smaller ranks are used to generate coarse features and rich contextual information, thereby preserving the generative capabilities of the original model. Meanwhile, lower timesteps receive a greater amount of information, facilitating the reproduction of concept fine details.
@@ 3.3 On the LoRA Orthogonality
Figure 4: Singular values of the B matrices for LoRA and Ortho-LoRA with $r=64$ after $800$ training iterations. The red line indicates the rank that encompasses 95% of the singular values total sum. The actual rank of the LoRA matrices is lower than full rank, particularly in the cross-attention layers, whereas the Ortho-LoRA matrices maintain full rank.
Masking matrix columns raises the issue of orthogonality, as non-orthogonal columns may have linear dependencies, limiting effective information exclusion. Our analysis of LoRA weights in diffusion model personalization shows that their effective rank is often much smaller than the specified rank (see Figure 4(a)), indicating redundancy among rank-1 components. This reduces the effectiveness of the Vanilla T-LoRA masking strategy and limits the model’s capacity utilization.
Forcing orthogonality in $A$ and $B$ matrices might resolve this issue. One approach involves using SVD-like architecture and regularization as in AdaLoRA adalora :
$\tilde{W}=W+BSA,\quad L_{reg}=\lambda_{reg}(\|AA^{T}-I\|^{2}_{F}+\|B^{T}B-I\|^{2}_{F})$ (4)
where $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ are initialized normally and $S\in diag(\mathbb{R}^{r})$ initialized with zeros. AdaLoRA adalora showed that scheme requires around 10,000 iterations to achieve orthogonality, far exceeding the 2,000 iterations typical for diffusion model personalization. Our experiments confirm that conventional initialization fails to achieve orthogonality within this limited schedule, making orthogonal initialization essential. Additionally, initializing $S$ matrices with zeros significantly slows training (see Section 4.1 for details).
To sum up we have the following restrictions in this task: (1) matrices $A$ and $B$ should be orthogonal from the beggining (2) matrix $S$ can not be initialized with zeros. To overcome these limitations, we employ the following LoRA trick, that allows to revisit LoRA weight initialization:
@@
The equation above allows us to bypass LoRA weights initialization requirement ($B=0$) and unlocks possibility to initialize LoRA with any type of weights. Specifically, as illustrated in Figure 2(c), we initialize matrices $A$ and $B$ with factor matrices derived from the Singular Value Decomposition (SVD) golub1971svd to enforce orthogonality, such that $A_{init}=V^{T}_{r}$, $B_{init}=U_{r},S_{init}=S_{r}$. We refer to this initialization method as Ortho-LoRA due to its inherent orthogonality properties.
The choice of SVD for initialization is crucial. We explore six variants of Ortho-LoRA initialization: using top, middle, and bottom singular components from the decomposed original weights $W\in\mathbb{R}^{n\times m}$, and a randomly initialized matrix $R\in\mathbb{R}^{n\times m}$. If we initialize the matrices with the top SVD components of $W$, the architecture simplifies to PISSA pissa . However, we argue that this initialization is not optimal for diffusion model customization. Our findings indicate that using high singular values correlate with overfitting, while using low singular values can slows down training (Figure 5). We show that using the last SVD components of a random matrix $R\in\mathbb{R}^{n\times m}$ works best for diffusion model customization (for more details, please refer to 4.1).
Finally, Figure 4(b) shows that in contrast to LoRA, Ortho-LoRA matrices retain full rank even after training, eliminating the need for additional orthogonal regularization.

### v2

@@ Abstract
While diffusion model fine-tuning offers a powerful approach for customizing pre-trained models to generate specific objects, it frequently suffers from overfitting when training samples are limited, compromising both generalization capability and output diversity. This paper tackles the challenging yet most impactful task of adapting a diffusion model using just a single concept image, as single-image customization holds the greatest practical potential. We introduce T-LoRA, a Timestep-Dependent Low-Rank Adaptation framework specifically designed for diffusion model personalization. We show that higher diffusion timesteps are more prone to overfitting than lower ones, necessitating a timestep-sensitive fine-tuning strategy. T-LoRA incorporates two key innovations: (1) a dynamic fine-tuning strategy that adjusts rank-constrained updates based on diffusion timesteps, and (2) a weight parametrization technique that ensures independence between adapter components through orthogonal initialization. Extensive experiments on SD-XL and FLUX-1.dev show that T-LoRA and its individual components outperform standard LoRA and other diffusion model personalization techniques, achieving a superior balance between concept fidelity and text alignment.
Project Pagehttps://controlgenai.github.io/T-LoRA/
@@ 1 Introduction
Recent advances in personalizing large-scale text-to-image diffusion models (Ruiz et al. 2023; Gal et al. 2022; Kumari et al. 2023) have revolutionized content creation, enabling pre-trained models with the ability to generate highly specific and customized outputs, such as particular objects, styles, or domains. The primary objectives of such customization are twofold: (1) ensuring high-quality preservation of the target concept, and (2) achieving precise alignment between the generated image and the input text.
Fine-tuning based customization approaches are effective at producing high fidelity concept samples (Ruiz et al. 2023; Kumari et al. 2023). However, they encounter limitations due to the restricted dataset size, which hinder generalization and cause artifacts like background elements and pose information to ”leak” into outputs, reducing the model’s flexibility and creativity. Compared to standard fine-tuning, lightweight Low-Rank Adaptater (LoRA) (Hu et al. 2021) offers significant advantages (Biderman et al. 2024; Ryu et al. 2024): it significantly reduces the number of trainable parameters, making it suitable for resource-constrained settings, and it is less prone to overfitting, thereby better preserving the original model’s generative capabilities.
Customizing diffusion models with a single concept image is particularly challenging due to the high risk of overfitting, even for lightweight methods (Qiu et al. 2023; Hu et al. 2021; Han et al. 2023) (Figure LABEL:fig:visual_abs). However, this task holds significant practical value, as users often lack multiple images of their concepts with varied backgrounds, making single-image customization a key focus of our research.
@@
Based on our analysis, we introduce T-LoRA, a Timestep-Dependent Low-Rank Adaptation framework for diffusion model customization. T-LoRA prioritizes training capacity for less noisy timesteps while reducing signals for noisier ones using a time-dependent masking strategy, Vanilla T-LoRA, which restricts higher-rank LoRA components during noisier timesteps. Our further analysis reveals that standard LoRA adapters often exhibit an effective rank that is significantly smaller than the original rank hyperparameter. This property could limit the effectiveness of Vanilla T-LoRA, as masked and unmasked adapters may possess similar expressive power. To address this, we introduce Ortho-LoRA, a novel LoRA initialization technique ensuring orthogonality between adapter components, thereby explicitly separating information flows across different timesteps. Extensive experiments demonstrate the effectiveness of T-LoRA and its components (Vanilla T-LoRA and Ortho-LoRA). Our proposed framework significantly outperforms existing lightweight fine-tuning approaches in single-image diffusion model personalization tasks in both metrics and user study. These results highlight the potential of integrating time-dependent regularization and orthogonality into future diffusion model adaptation frameworks.
@@
  • We propose T-LoRA, a Low-Rank Adaptation framework that mitigates overfitting in diffusion model personalization through a rank masking strategy, balancing training capacity and applying stronger regularization at higher timesteps.
  • We explore the concept of effective rank in LoRA matrices and propose Ortho-LoRA, a novel orthogonal weight initialization method, that enhances effective rank utilization and improves information flow separation across timesteps, boosting T-LoRA performance.
@@ 2 Background
Diffusion models (Song et al. 2020; Ho et al. 2020) are probabilistic generative models using neural networks to approximate data distributions by iteratively denoising Gaussian-sampled variables. We focus on text-to-image diffusion models $\varepsilon_{\theta}$, which generate images $x$ from text prompt $P$. These models use a text encoder $E_{T}$ to extract text embeddings $p=E_{T}(P)$. Latent diffusion models (Rombach et al. 2022) encode images into latent representations $z=E(x)$ via an encoder $E$ and decode them with a decoder $D$, ensuring $x\approx D(z)$. The diffusion model $\varepsilon_{\theta}$ is trained to predict noise:
$\min_{\theta}\mathbb{E}_{p,t,z,\varepsilon}\left[\left\|\varepsilon-\varepsilon_{\theta}(t,z_{t},p)\right\|_{2}^{2}\right],$ (1)
@@
where $\varepsilon\sim N(0,I)$, $t$ is the diffusion timestep, and $z_{t}$ is a noisy latent code obtained via the forward process $z_{t}=\text{ForwardDiffusion}(t,z_{0},\varepsilon)$. During inference, random noise $z_{T}\sim N(0,I)$ is iteratively denoised to recover $z_{0}$.
@@
Low-Rank Adaptation (LoRA) (Hu et al. 2021) efficiently fine-tunes large pre-trained models by updating the weight matrix $W$ as $\tilde{W}=W+BA$, where $A\in\mathbb{R}^{r\times m}$, $B\in\mathbb{R}^{n\times r}$, $r$ is the rank, and $(n,m)$ is the dimensionality of $W$. To preserve the pre-trained model’s behavior initially, $A$ is normally initialized, and $B$ is set to zero.
@@
Diffusion Model Customization often involves fine-tuning, where model weights are adjusted to generate specific user-defined concepts from a limited set of concept images. To associate a new concept with a special text token $V^{*}$, $\varepsilon_{\theta}$ is fine-tuned on a small set of concept images $\mathbb{C}=\{x\}_{i=1}^{N}$ using the following objective:
$\min_{\theta}\mathbb{E}_{p,t,z=\mathcal{E}(x),x\in\mathbb{C},\varepsilon}\left[\left\|\varepsilon-\varepsilon_{\theta}(t,z_{t},p)\right\|_{2}^{2}\right],$ (2)
where $p=E_{T}(P)$ represents the text embeddings for the prompt $P=$ ”a photo of a V*”. Fine-tuning enhancements include pseudo-token optimization (Gal et al. 2022), diffusion model fine-tuning (Ruiz et al. 2023; Kumari et al. 2023), and lightweight parameterization (Han et al. 2023; Qiu et al. 2023) to reduce costs and mitigate overfitting. LoRA (Hu et al. 2021), with its lightweight design, high concept fidelity, and strong prompt alignment, serves as a strong baseline for subject-driven generation and a key component in both fine-tuning-based (Arar et al. 2024) and encoder-based (Li et al. 2024) personalization techniques.
@@
Overfitting Problem Existing methods often overfit to position and background, especially with limited concept images. To address this, various techniques like image masking (Wei et al. 2023; Huang et al. 2024), prompt augmentation (Sohn et al. 2023; Garibi et al. 2025), regularization (Kumari et al. 2023; Wu et al. 2025; Pang et al. 2024), advanced attention (Huang et al. 2024; Nam et al. 2024), and sampling (Zhou et al. 2023; Soboleva et al. 2025) have been explored. In contrast, we identify the root cause as fine-tuning during noisiest timesteps, which reinforces background elements and limits flexibility. To address this, we propose reducing concept signals during noisy timesteps and adapting LoRA (Hu et al. 2021) to control concept injection across timesteps, enhancing generalization and diversity while mitigating overfitting.
Figure 2: Motivational Experiment. The figure shows fine-tuning the SD-XL model with LoRA over fixed timestep intervals. Focusing on the noisiest timesteps causes rapid overfitting, affecting positioning and backgrounds. Shifting to earlier timesteps improves text alignment and generation flexibility. However, completely excluding the noisiest timesteps ($t\in[800;1000]$) is infeasible, as they are essential for maintaining concept fidelity.
@@ 3 Method
Motivation Previous studies have shown that different timesteps in diffusion models play distinct roles throughout the generation process (Choi et al. 2022; Deja et al. 2022; Li et al. 2023). For example, the authors of (Choi et al. 2022) categorized the behavior of diffusion timesteps into three main stages: high timesteps ($t\in[800;1000]$) concentrate on forming coarse features, middle timesteps ($t\in[500;800]$) produce perceptually rich content, and the lower timesteps ($t\in[0;500]$) focus on removing residual noise. Additionally, works (Chang et al. 2023; Gao et al. 2023) show that high timesteps contribute to image diversity, as inadequate representation of the prompt’s context during this stage makes it less likely to be restored in subsequent timesteps. Leveraging these insights, a growing body of work has introduced techniques to optimize diffusion model training. For instance, (Wang et al. 2025) proposed a time-dependent loss weighting strategy, while (Kim et al. 2025) and (Zheng et al. 2024) designed adaptive timestep sampling methods to improve the overall generation quality and diversity.
To investigate the role of different timesteps we fine-tuned the SD-XL model with LoRA over fixed timestep intervals (see Figure 2). The results show that fine-tuning at higher timesteps $t\in[800;1000]$ leads to rapid overfitting, causing memorization of poses and backgrounds, which limits image diversity and prompt alignment, though these timesteps are crucial for defining shape and proportions. In the middle timesteps $t\in[500;800]$, the generated context became richer, and the model better reproduce fine concept details. However, we lose information related to the overall shape. For example, the boot in Figure 2 retains fittings that correspond to the original concept, but the shoe is shorter. Finally, fine-tuning with lower timesteps $t\in[0;500]$ demonstrated the best alignment with text prompts and yielded the richest generation. However, this approach struggled to accurately reproduce the intended concept, losing both shape and fine details of the object.
@@
Figure 3: Comparison of training methods: LoRA, the proposed Vanilla T-LoRA, and T-LoRA schemes.
@@ 3.1 Vanilla T-LoRA
To tackle the aforementioned challenge, we propose a timestep-dependent fine-tuning strategy for diffusion models. Our method dynamically adjusts the ranks of LoRA adapters based on the diffusion timestep, allocating fewer parameters at higher timesteps and more at lower ones. This approach, called Vanilla T-LoRA, incorporates a masking mechanism (see Figure 3(b)):
$\centering\begin{split}&\tilde{W}_{t}=W+B_{t}A_{t}=W+BM_{t}A,\ A\in\mathbb{R}^{r\times m},B\in\mathbb{R}^{n\times r}\\ &M_{t}=M_{r(t)}=\mathrm{diag}(\underbrace{1,1,\dots,1}_{r(t)},\underbrace{0,0,\dots,0}_{r-r(t)})\in\mathbb{R}^{r\times r}\end{split}\@add@centering$ (3)
We define $r(t)$ as a linear function inversely proportional to timesteps: $r(t)=\lfloor(r-r_{\text{min}})\cdot(T-t)/T\rfloor+r_{\text{min}}$, where $r_{\text{min}}$ is a hyperparameter. This rank-masking strategy dynamically controls information injection across timesteps during training and inference. Higher timesteps use smaller ranks to preserve generative capabilities by focusing on coarse features and context, while lower timesteps receive more information to capture fine concept details.
@@ 3.2 On the LoRA Orthogonality
Figure 4: Singular values of B matrices for LoRA and Ortho-LoRA with $r=64$ after 800 training steps. The red line marks the rank capturing 95% of the total singular value sum. LoRA matrices are effectively low-rank – especially in cross-attention – while Ortho-LoRA maintains full rank.
Linear dependence among the columns of LoRA matrices can compromise the effectiveness of the Vanilla T-LoRA masking strategy, limiting the ability to exclude information effectively. Our analysis of LoRA weights in diffusion model personalization demonstrates that their effective rank is frequently much smaller than the specified rank (Figure 4(a)), demonstrating linear dependency between the matrix columns. Enforcing orthogonality in $A$ and $B$ matrices could address this, using an SVD-like architecture and regularization, as in AdaLoRA (Zhang et al. 2023):
$\centering\begin{split}&\tilde{W}=W+BSA,\\ &L_{reg}=\lambda_{reg}(\|AA^{T}-I\|^{2}_{F}+\|B^{T}B-I\|^{2}_{F})\end{split}\@add@centering$ (4)
where $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ are normally initialized, and $S\in diag(\mathbb{R}^{r})$ is zero-initialized. AdaLoRA demonstrated that this setup requires approximately 10,000 iterations to achieve orthogonality, which is significantly higher than the 1,000-2,000 iterations typically needed for diffusion model customization. As a result, this regularization approach is not suitable for personalization tasks, making orthogonal initialization essential. Furthermore, initializing the $S$ matrices with zeros considerably slows down the training process (see Section 4.1 for more details).
In summary, the task requires: (1) $A$ and $B$ to be orthogonal from the start, and (2) $S$ not to be zero-initialized. To address these, we use a LoRA trick to revise LoRA weight initialization:
@@
This removes the need for zero initialization ($B=0$), enabling arbitrary weight initialization. As shown in Figure 3(c), we initialize $A$ and $B$ using SVD (Golub and Reinsch 1971) factor matrices to enforce orthogonality: $A_{init}=V^{T}_{r}$, $B_{init}=U_{r},S_{init}=S_{r}$, and $S_{init}=S_{r}$. We term this approach Ortho-LoRA due to its orthogonal structure.
The choice of SVD for initialization is crucial. We examine six Ortho-LoRA variants: using top, middle, and bottom singular components of original weights $W\in\mathbb{R}^{n\times m}$, and of a random matrix $R\in\mathbb{R}^{n\times m}$. Top-component initialization from original weights reduces to PISSA (Meng et al. 2024), but we find it suboptimal for diffusion model customization. Initializing with the last SVD components of random matrix $R$ yields optimal results (see Section 4.1). Figure 4(b) illustrates that, in contrast to LoRA, Ortho-LoRA maintains full rank throughout the entire training process without requiring any orthogonal regularization.